OpenStudy (anonymous):

solve \[n\cdot \log_2 =10^9\] for n ?

OpenStudy (anonymous):

log2 of what?

OpenStudy (anonymous):

my bad \[ n \cdot \log_2 n =10^9\]

OpenStudy (anonymous):

No problem. So that gives us log2n^n=10^9.

OpenStudy (anonymous):

So far, the clues we have about n is that it needs to be a multiple of 2.

OpenStudy (anonymous):

right, and also you can right it \[ \frac{ln (n^n)}{ln (2)}= 10^9\] \[ ln (n^n)= 10^9 \cdot ln (2)\]

OpenStudy (anonymous):

write*

OpenStudy (anonymous):

how to solve the LHS?

OpenStudy (anonymous):

Let's form an equation to solve this, shall we? \[\frac {2^{n}}{2} \times n = 10^{9}\]

OpenStudy (anonymous):

I am not sure, \[n\cdot log_2(n)\]=\[e^{(\frac{n*ln(n)}{ln(2)})}\]=\[n^{(\frac{n}{ln(2)})}\]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Hero , @iambatman

OpenStudy (anonymous):

@zepdrix "NOON"

OpenStudy (anonymous):

@swaschan \[\frac{n\cdot2^n}{2} \neq n\cdot log_2(n)\]

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

@dumbcow

OpenStudy (zzr0ck3r):

there are no basic ways to solve this. We need something called the product log, http://www.wolframalpha.com/input/?i=x*log_2%28x%29%3D10%5E9

OpenStudy (jhannybean):

@wio how do these work?

OpenStudy (anonymous):

Raise both sides to the power of \(2\), and:\[ n^n=2^{10^9} = (1024)^{10^8} \]We want the base and exponent to have the same number of digits.

OpenStudy (anonymous):

Raise both sides to the power of \(2\), and:\[ n^n=2^{10^9} = (1024)^{10^8} \]We want the base and exponent to have the same number of digits. Base has \(4\), while exponent has \(8\). If we divide exponent in half, which would square the base, we would have the base having \(8\) digits and the exponent having \(7\). So \(n\approx 1024^2\) is a good start. You would want to use newton's method to get more accurate answers. Either way it seems like \(10^7<n<10^8\).

OpenStudy (perl):

2^(10^9) =/= (2^10)^9

OpenStudy (perl):

n^ n = 2^(10^9) n^n = 2^ 1000000000

OpenStudy (perl):

n = e^(W(1000000000 log(2))) ~ 3.96201*10^7

OpenStudy (perl):

W(z) is the product log function

OpenStudy (anonymous):

\[A_{n \times 1}, B_{n \times n} , C_{n \times 1}\] are three matrix, proof that \[A^TBC=\sum_{i=1}^{n}B_{ij}[AC^T]_{ji}\]